Calculate the molarity of a solution prepared by dissolving 24.5 g of KF in enough water to produce 376 mL of solution
1
Expert's answer
2020-02-05T06:13:58-0500

Molarity is defined as a number of moles of solute in 1 L of solution. As a result, it csn be calculated using the following formula:

M = n / V = m / (Mr × V)

where M – molarity, n – number of moles, m -mass, Mr – molecular weight, V – volume.

As Mr(KF) = 58 g/mol and V = 376mL = 0.376 L:

M (KF) = 24.5 g / (58 g/mol × 0.376 L) = 1.1 mol/L = 1.1 M


Answer: 1.1 M